OpenStudy (anonymous):

Simplify & Solve: 3y+5=y-25

OpenStudy (anonymous):

3y+5=y-25 3y=y-30 2y=-30 y=-15

OpenStudy (anonymous):

For these ones just try to arrange the ys on one side and the numbers on the other. so take -y from both sides and -5 from both sides and than divide by two